### v1

@@ Appendix A Generating a synthetic MPEC benchmark
To generate a problem test set, we consider 30 objective functions. Two of them correspond to positive definite and indefinite quadratic objectives, of the form: $f(x)=(x_{0},x_{1})^{\top}H(x_{0},x_{1})+g^{\top}(x_{0},x_{1})$, where $H=0.5(\hat{H}^{\top}+\hat{H})$, $\hat{H}\sim\mathcal{U}_{(p+m)\times(p+m)}(0,1,s)$ and vector $g\sim\mathcal{U}_{p+m}(-10,10)$. Only in these two cases, we do not make the equality and inequality constraints nonlinear. For the other 28 functions, we use the following nonlinear test functions from Andrei2008 : Fletcher, Himmelblau, McCormick, Powell, Trigonometric, Rosenbrock, Raydan1, Raydan2, Diagonal3, Diagonal4, Diagonal5, Extended_Trigiaonal, Three_Exponential_Terms, Generalized_PSC1, Extended_PSC1, Fletcvb3, Bdqrtic, Tridia, EG2, Edensch, Indef, Cube, Bdexp, Genhumps, Arwhead, Quartc, Cosine, Sine. For every test function, we generate $7$ problem instances with different sizes, resulting in a total of $210$ MPECs. The dimensions of $p$ are unique random integers from the set $\{10,\ldots,300\}$, $p=m$ and $n_{\mathrm{ineq}}~{}\mathrm{round}(\mathcal{U}_{1}(0.1p,2p))$. The remaining parameters we use in Algorithm 4 are $\alpha=10$, $s^{\mathrm{eq}}=0.5$, $s^{\mathrm{ineq}}=0.5$, and $s^{d}=0.5$.

### v2

@@ Appendix D Generating a synthetic MPEC benchmark
To generate a test problem set, we consider 25 objective functions and create three instances of different sizes for each problem, resulting in a total of 75 problems. The the nonlinear objective functions , we use the following nonlinear test functions from Gould2013 : Fletcher, McCormick, Powell, Rosenbrock, Raydan1, Raydan2, Diagonal4, Diagonal5, Extended_Trigiaonal, Three_Exponential_Terms, Generalized_PSC1, Fletcvb3, Bdqrtic, Tridia, EG2, Edensch, Indef, Cube, Bdexp, Genhumps, Arwhead, Quartc, Cosine, NCVXQP6, DIXCHLNV. The dimensions of $p$ are unique random integers from the set $\{100,\ldots,2100\}$, $p=m$ and $n_{\mathrm{ineq}}~\mathrm{round}(\mathcal{U}_{1}(0.5p,2p))$. The remaining parameters we use in Algorithm 4 are $\alpha=10$, $s^{\mathrm{eq}}=0.01$, $s^{\mathrm{ineq}}=0.25$, and $s^{d}=0.25$.